find the length of each side of a triangle whose sides have midpoints at (3,2,3)(-1,1,5) and (0,3,4).
Answer by psbhowmick(878)

Let the coordinates of the vertices be (x1,y1,z1), (x2,y2,z2) and (x3,y3,z3).

Then use the following equations:
(x1+x2)/2=3
(x2+x3)/2=-1
(x3+x1)/2=0
Etc.

Solve these 9 equations and get the values of 9 unknowns.

Cheers!